OpenStudy (harleyquinn):

How do I write this in standard form? y=6x-11

OpenStudy (anonymous):

ax+by= c is standard form

OpenStudy (harleyquinn):

You sure? I thought it was A(sq)+B(sq)=C(sq)...Oh wait is that Pythagorean theorem?

OpenStudy (anonymous):

Yes thats the pythagorean theorm for determining sides of right triangles

OpenStudy (harleyquinn):

Ah. Okay. So how do I put y=6x-11 into standard form?

OpenStudy (anonymous):

it 1y and 6x and the constant or number is c see if you can figure it out with that information

OpenStudy (harleyquinn):

I honestly don't know.

OpenStudy (harleyquinn):

-6x+y=-11?

OpenStudy (anonymous):

Thats close but the beginning can't be negative so if you multiply everything by -1 it would be 6x-y=11

OpenStudy (harleyquinn):

the beginning can't be negative?

OpenStudy (harleyquinn):

Ever?

OpenStudy (anonymous):

in every class ive been in they have told me that.

OpenStudy (anonymous):

I haven't seen an exception

OpenStudy (harleyquinn):

Okay thanks.

OpenStudy (anonymous):

No problem